# Free Record-Level Privacy Risk Evaluation Through Artifact-Based Methods

### v1

@@ 3.1.2 Membership Inference Attacks
Throughout this work, we focus on identifying the most vulnerable samples by fixing a low FPR threshold $\alpha$ (typically $\alpha=10^{-3}$) and considering samples that are identified as members at this strict threshold by SOTA shadow model-based MIAs. We use the terms “point” and “sample” interchangeably. For a given $\alpha$, we select the threshold $\tau_{\alpha}$ such that $\text{FPR}(\tau_{\alpha})=\alpha$.
Definition 3.1 (Vulnerable points).
3.1.3 Likelihood Ratio Attack (LiRA)
LiRA [7] frames membership inference as a likelihood ratio test between two hypotheses. For a target point $(x,y)$, it compares the likelihood of observing the model’s loss under the hypothesis that $(x,y)$ was a training set member versus a non-member. Specifically, it estimates:
$A_{\text{LiRA}}(x,y,\theta)=\log\frac{p(\ell(f_{\theta}(x),y)|(x,y)\in D_{\text{train}})}{p(\ell(f_{\theta}(x),y)|(x,y)\not\in D_{\text{train}})}$ (9)
To estimate these likelihood distributions, LiRA trains $N$ pairs of shadow models $\{(M_{i}^{+},M_{i}^{-})\}_{i=1}^{N}$, where $M_{i}^{+}$ includes $(x,y)$ in its training set and $M_{i}^{-}$ excludes it. The shadow models provide samples from the loss distributions conditioned on membership and non-membership, allowing estimation of the likelihood ratio.
3.1.4 Attack R
Attack R [40] frames membership inference as a sample-dependent test by comparing the target model’s loss against reference models trained without the target point. For a target point $(x,y)$, it computes:
$A_{\text{Attack R}}(x,y,\theta)=\text{Pr}[\ell(f_{\theta}(x),y)\leq c_{\alpha}(x,y)]$ (10)
where $c_{\alpha}(x,y)$ is a threshold that depends on the specific target point $(x,y)$. To estimate this threshold, Attack R trains $N$ reference models $\{M_{i}\}_{i=1}^{N}$ on randomly sampled datasets that exclude $(x,y)$. The threshold $c_{\alpha}(x,y)$ is set as the $\alpha$-percentile of the empirical loss distribution $\{\ell(f_{M_{i}}(x),y)\}_{i=1}^{N}$ to achieve a target false positive rate $\alpha$.
Contrary to LiRA, Attack R only trains reference models without the target point (rather than pairs with/without), making it more computationally efficient.
@@ 4 Method
(a) Loss traces of 3 samples from the “frog” class in CIFAR-10 representing different example hardness.
(b) Average loss traces of 2 groups of 100 points representing different example hardness and the average proportion of vulnerable points per group, compared to the average test loss.
@@ 4.1 Intuition
The main challenge of running loss-based MIAs stems from the fact that the final loss for a given sample is influenced by two independent factors: example hardness and memorization [7]. For instance, extremely low loss values can either represent a very easy-to-classify non-member example or a well-memorized member. Similarly, a relatively high loss value can represent either a hard-to-classify, partially memorized member, or an average non-member. Identifying whether a given sample is easy- or hard-to-classify is therefore a key component of a strong loss-based MIA.
@@
MIAs relying on shadow models [7, 40] account for example hardness by training a number of models on datasets without the target sample. The resulting loss distributions of the target sample are then measured, and used to inform the attack.
@@
We here posit that example hardness might also be inferred, for free, from training artifacts such as the loss trace over training. The intuition is that easy-to-fit samples would exhibit an early drop in the training loss and quickly reach a small loss value. Meanwhile, hard-to-classify yet memorized samples would take more epochs to converge to a final, and equally small, loss value. A similar logic applies to samples with moderately high losses at the end of training. If they are very hard-to-classify samples exhibiting even higher losses on non-member shadow models, they could still be correctly identified by an MIA. Here, a steep drop in loss during training (even if not reaching zero) would signal that a sample is partially memorized.
@@
Fig. 2(a) shows loss traces on three samples from the CIFAR-10 “frog” class, collected during training (all samples were part of the training dataset). These traces represent distinct types of loss behavior. The purple sample is a typical sample for the frog class: it is easily learned by the model and the train loss quickly converges to small loss values. The olive sample, on the other hand, is an outlier frog sample. It features unique characteristics compared to other samples from the same class, notably the color. Here we can see that the loss is quite noisy and remains fairly high until approximately epoch 70 when it gets memorized. Finally, the cyan sample is an even harder-to-predict sample from the frog class. The early loss (around epoch 10) for this sample is significantly higher than for the purple example. The loss then decreases, but does not reach the low values exhibited by the purple and olive samples by the end of training.
@@
Following this intuition, we further verify our hypothesis by grouping samples with similar loss traces together, and measuring the proportion of samples considered vulnerable by LiRA [7] at $\alpha=10^{-3}$ (see Def. 3.1) in these groups. In particular, on Figure 2(b) we consider average loss traces for two groups with 100 samples each, alongside the average train loss. Both groups are selected to have relatively high loss at an early sample in the training (epoch 11, vertical dashed line on the graph) making them more likely to be outliers and at-risk samples. One group then converges to near-zero training loss, representing easy-to-fit outliers. The final loss for the final group remains relatively high, thus representing hard-to-fit outliers. Borrowing the nomenclature from Carlini et al. [7], we expect a) outliers to be more vulnerable to MIAs and b) easy-to-fit samples to be more vulnerable than hard-to-fit samples. Figure 2(b) indeed confirms that vulnerable samples make up 29% of the easy-to-fit outlier group compared to 12% for hard-to-fit outliers. When comparing to the average baseline of 5%, we confirm that samples with high loss at the early stage of training, outliers, are more vulnerable to MIAs and that, among outliers, easy-to-fit samples (defined by the low final loss) are more vulnerable than hard-to-fit ones. Our method aims to utilize this information to identify at-risk samples.
4.2 Computing per-sample losses
In most cases, computing the per-sample loss during training incurs little overhead in addition to normal training. For instance, we consider the typical PyTorch training loop below, where outputs and targets represent batched data.
criterion = nn.CrossEntropyLoss()
loss = criterion(outputs, targets)
# here the loss is averaged out and
# does not have a batch dimension
loss.backward()
Here the loss reduction is set to mean by default and the loss is averaged over a batch before being returned from the criterion call. Instead, we can easily perform the aggregation on the client side, storing the per-sample losses before averaging.
criterion = nn.CrossEntropyLoss(
reduction="none"
loss = criterion(outputs, targets)
# here the loss is per-sample loss
# and does have a batch dimension
saved_losses.append(
loss.detach()
# take the mean before backward pass
loss.mean().backward()
We note, however, that this approach is not applicable if augmentations are used. In this case tracking per-epoch losses incurs the cost of one forward pass on the full dataset on every epoch.
Memorization is indeed a property of a specific sample: the model’s behaviour on memorized samples and their augmented versions might typically be very different. As such, when augmentations are applied during training, losses computed during training would be losses on augmented samples, not the original ones. In this scenario we suggest recording the losses of the non-augmented training samples in a separate forward pass after each epoch. While slightly more costly, this approach remains computationally much cheaper than training even one shadow model, as a forward pass is typically several times cheaper than the backward pass.
@@ 4.3 Loss trace aggregation
Note that, where relevant, early epoch $s^{*}$ is defined as an epoch at which the average test loss $\mathcal{L}_{\text{test}}^{s^{*}}$ has decreased significantly compared to the initial $\mathcal{L}_{\text{test}}^{0}$, but the train/test gap ($\mathcal{L}_{\text{test}}^{s^{*}}-\mathcal{L}_{\text{train}}^{s^{*}}$) remains low. Intuitively, we consider epoch $s^{*}$ as a point where the model has gained some generalization ability, but has not significantly memorized any samples.
@@
Final loss
This represents the final model loss at epoch $S$.
$\phi(x,y)=\ell_{S}(x,y)$ (11)
Mean loss
@@
Loss delta
@@
Smooth loss delta
Intuitively, this represents the same property as loss delta, but the values are averaged over multiple epochs (for a certain value of $\delta$) for robustness.
$\phi(x,y)=\frac{1}{2\delta}\sum_{s=s^{*}-\delta}^{s^{*}+\delta}\ell_{s}(x,y)-\frac{1}{2\delta}\sum_{s=S-2\delta}^{S}\ell_{s}(x,y)$ (14)
Normalized loss delta
It measures the fraction of a sample’s early-stage loss that was eliminated by the end of training, capturing the relative (rather than absolute) improvement in the model’s performance on that sample.
$\phi(x,y)=\frac{\ell_{s^{*}}(x,y)-\ell_{S}(x,y)}{\ell_{s^{*}}(x,y)}$ (15)
@@ 5.1 Experimental setup
The code to reproduce the experiments is available at https://github.com/computationalprivacy/loss_traces.
@@ 5.1.1 Target model
Adopting the evaluation setup from the literature [7], we train Wide ResNet28-2 [43] models on the CIFAR-10 dataset [19] for 100 epochs, reaching a test accuracy of $\sim$86%. We train using a stochastic gradient descent optimizer, with a momentum of 0.9 and weight decay set to 0.0001. We use a cosine annealing learning rate scheduler with initial value of 0.1. We use a batch size of 256. During training, we augment the images via random horizontal flips, and random crops with four pixels of padding, following the augmentation setup used for training by Carlini et al. [7].
@@
Where appropriate, we set the early stage epoch $s^{*}$ (see Sec. 4.3) at 11. At this stage the average test loss $\mathcal{L}_{\text{test}}^{s^{*}}$ has decreased by $\sim 75$%, while the train/test gap $\mathcal{L}_{\text{test}}^{s^{*}}-\mathcal{L}_{\text{train}}^{s^{*}}$ remains low at $\sim 2$%.
@@ 5.1.2 Membership Inference Attacks
We evaluate three distinct membership inference attacks: LiRA, Attack R, and LOSS. We follow the setup proposed in Carlini et al. [7], instantiating online LiRA with 256 shadow models (128 member and 128 non-member models respectively). We then also instantiate Attack R with 128 non-member shadow models. All shadow models are trained using the same training procedure described for the target model in Sec. 5.1.1, with each shadow model trained on a different member/non-member split. For LiRA, we use 2 queries per sample through horizontal flipping, matching the “mirror augmentation” setup from Carlini et al. [7].
Figure 3: ROC curves for LiRA, Attack R, and LOSS membership inference attacks on CIFAR-10 against Wide ResNet28-2 (log scale).
MIA AUC TPR@FPR=$10^{-3}$
LiRA [7] $0.71$ $0.049$
Attack R [40] $0.64$ $0.013$
LOSS [41] $0.57$ $0.001$
Table 1: AUC and TPR@FPR=$10^{-3}$ scores for LiRA, Attack R and LOSS MIAs with our training setup.
Table 1 shows the performance of MIAs we consider in our training setup. We evaluate them on a random balanced split of members and non-members from the CIFAR-10 train split. As expected, SOTA shadow model-based attacks (LiRA, Attack R) significantly outperform LOSS, with LiRA showing the best performance. Note however that though the LOSS attack shows non-trivial AUC=$0.57$, it fails to confidently identify members - as measured by TPR@FPR=$10^{-3}$ being at random chance. Additionally, Fig. 3 shows full AUC curves for the three attacks.
5.1.3 Evaluation metric
We evaluate our vulnerability detection methods using Precision@k and Recall@k, which measures what fraction of the k points with highest vulnerability scores are actually vulnerable according to a reference MIA (LiRA). For a vulnerability scoring function $\phi$ and a set of vulnerable points $V_{\alpha}$ (for a given FPR threshold $\alpha$ as defined in Definition 3.1), we define Precision@k as:
[…]
where $\text{top}_{k}(\phi)$ represents the k points with the highest vulnerability scores according to $\phi$. In other words, if we sort all training points by their vulnerability score $\phi(x,y)$ in descending order and take the first k points, Precision@k tells us what fraction of these points are actually vulnerable according to our reference MIA.
Similarly, we define Recall@k as:
[…]
Recall here measures the fraction of all vulnerable points we successfully identified within our top k predictions.
The results we report are averaged across 10 models trained with varying randomization seed. Throughout the paper, we use LiRA as the reference MIA.
5.2 Cross-MIA correlation
Figure 4: Confusion matrix comparing the predictions of vulnerable and non-vulnerable points identified by LiRA and Attack R where we consider the top 1000 points to be vulnerable.
Before evaluating our method’s effectiveness at identifying vulnerable points, it’s important to understand the inherent challenges of this task.
First, vulnerable points are rare: at a strict false positive rate of $10^{-3}$, LiRA, for instance, identifies only 1,153 samples to be at-risk, slightly below 5% of our training set of 25,000 samples. This means that finding vulnerable points is a significantly more challenging task than for instance the standard member vs. non-member classification. Indeed, even small absolute errors can lead to large relative performance drops when dealing with such a small target set and a random guess baseline for this task would be below 5%.
Furthermore, we study the level of agreement between SOTA methods on which points are the most at-risk and show, somewhat surprisingly, that there is only a partial agreement on which points are most vulnerable. Figure 4 shows the overlap between the top 1,000 most vulnerable points identified by each attack. Despite both LiRA and Attack R assuming the same attacker, being both shadow model-based and part of the same class of attacks, and achieving strong overall membership inference performance (AUC of 0.71 and 0.64 respectively, see Table 1), they only share 551 points among their top 1,000 predictions. We believe this to be both an important yet understudied aspect of MIAs, with potentially important practical consequences.
@@ 5.2 Cross-MIA correlation
Attack R [40] $0.62$ $0.13$ $0.60$ $0.37$ $0.57$ $0.59$
LOSS [41] $3\times 10^{-4}$ $7.2\times 10^{-5}$ $6.7\times 10^{-4}$ $3.7\times 10^{-4}$ $5.6\times 10^{-4}$ $5.4\times 10^{-4}$
Final loss $0.08$ $0.02$ $0.11$ $0.07$ $0.12$ $0.12$
Mean loss $0.36$ $0.07$ $0.38$ $0.24$ $0.36$ $0.37$
Loss delta $0.09$ $0.02$ $0.10$ $0.06$ $0.07$ $0.08$
Normalized loss delta $0.14$ $0.03$ $0.12$ $0.07$ $0.12$ $0.12$
LT-IQR (75%-25%) $0.61$ $0.13$ $0.49$ $0.30$ $0.42$ $0.43$
Table 2: Precision@k and Recall@k for various loss trace aggregation methods, reported at fixed FPR=$10^{-3}$ and varying values of $k$. We also report the performance of existing MIAs on the same task for the reference
Figure 5: Recall@k=1% (250 points) for identifying vulnerable points determined by LiRA attack at a variable FPR threshold (log scale).
5.3 LT-IQR empirical results
We report the precision of different vulnerability detection methods (Final loss, Attack R, us) and random guess at identifying points that are the most vulnerable, defined as LiRA at low FPR points.
Figure 1 shows that our method succeeds at identifying the most vulnerable points (a low FPR regime) according to LiRA, and does so as well as Attack-R, while being orders of magnitude computationally cheaper. Focusing on the FPR=$10^{-3}$ threshold, which corresponds to a highly selective attack setting, our best performing aggregation metric (LT-IQR (75%-25%)) achieves approximately 60% precision at $k=1\%$ of training data. This means that among the 250 training points our method identifies as most vulnerable, 61% are actually vulnerable according to LiRA at this strict false positive threshold. This, impressively, matches the performance of Attack R despite requiring no shadow model training compared to Attack R’s 128 shadow models. It also significantly outperforms simpler baselines like final loss (8% precision) and random guess.
Looking at the broader trend across FPR thresholds, we observe that LT-IQR’s performance advantage does persist. At stricter thresholds, LT-IQR slightly outperforms Attack R while strongly outperforming both final loss and random baselines. As we relax the FPR threshold (moving right on the graph), Attack R starts to show slightly higher precision compared to our method for medium-vulnerability points.
We then consider the recall of the predictions. Figure 5 evaluates the recall of different vulnerability detection methods across varying FPR thresholds, examining what fraction of truly vulnerable points each method successfully identifies within its top k=1% predictions. The results reveal several key patterns. First, at very strict thresholds ($\text{FPR}<10^{-3}$), LT-IQR identifies around 50% of vulnerable points in its highest-ranked predictions - meaning we capture half of these extremely vulnerable samples by predicting just 250 samples (1% of the training data). As we relax the FPR threshold, recall will then naturally decrease: while more points become classified as vulnerable at higher FPR thresholds, our method still examines only the same top $k=250$ predictions. The fact that LT-IQR maintains higher recall than Attack R until FPR=$10^{-3}$ (and closely matches it for higher FPR values), demonstrates its robustness as a sample vulnerability detection method. At more permissive thresholds ($\text{FPR}>0.01$), all methods, including random guessing, begin to achieve similar recall, reflecting the fact that the task becomes less discriminative when a large fraction of points are considered vulnerable.
Taken together, these results demonstrate that our simple, artifact-based LT-IQR method succeeds at the challenging task of identifying the points most vulnerable to membership inference attacks. The ability to achieve 60% precision at FPR=$10^{-3}$ is particularly notable given the difficulty of the task - we are identifying a small proportion of training points that can be confidently detected as members even under the most conservative attack settings that virtually eliminate false positives. The fact that we can match the performance of expensive SOTA shadow model-based approaches like Attack R, while using only training artifacts that are essentially free to collect, represents a significant advance in making privacy risk assessment more practical and accessible during model development. We believe that this high precision at strict FPR thresholds means model developers can reliably use LT-IQR to identify and address their models’ most severe privacy vulnerabilities during model development without the computational burden of training hundreds of shadow models.

### v3

@@ 4.2 Membership Inference Attacks
Throughout this paper we run the following MIAs:
  1. 1.
    LOSS [46]
  2. 2.
    Likelihood Ratio Attack (LiRA) [5]
  3. 3.
    Attack R [45]
  4. 4.
    Robust Membership Inference Attack (RMIA) [51]
4.3 Attack Risk Predictors
Previous works have proposed various metrics as indicators of potential vulnerability to privacy attacks. We select several key predictors commonly cited in the literature [22]:
Loss measures the discrepancy between a model’s prediction and the ground truth label and is used by many MIAs [45, 39, 26].
Confidence quantifies the margin between the model’s highest probability prediction and its second-highest prediction, and is also used as the basis of an MIA [5].
Input gradient norm captures the sensitivity of the model’s output with respect to small perturbations in the input features and can be used as a form of feature attribution [22].
Parameter gradient norm evaluates how changes to the model’s parameters influence its predictions. Gradient norms are leveraged by the GLiR Attack [20] and have been explored for links to vulnerability [22].
SHAP Values [29] leverages game theory to determine feature importance by measuring each input’s marginal contribution to the model’s predictions and has been considered in the context of understanding the privacy risks of model explanations [25].
@@ 5 Method
((a))
((b))
@@ 5.1 Intuition
The main challenge of running loss-based MIAs stems from the fact that the final loss for a given sample is influenced by two independent factors: example hardness and memorization [5]. For instance, extremely low loss values can either represent a very easy-to-classify example or a well-memorized one. Similarly, a relatively high loss value can represent either a hard-to-classify, partially memorized example, or an average non-memorized example.
@@
MIAs relying on shadow models [5, 45] aim to account for example hardness by training a number of models on datasets without the target sample. The resulting loss distributions of the target sample are then measured, and used to inform the attack.
@@
We here posit that example hardness might also be inferred, for free, from training artifacts such as the loss trace over training, and then used to evaluate the corresponding privacy risk. The intuition is that easy-to-fit samples would exhibit an early drop in the training loss and quickly reach a small loss value. Meanwhile, hard-to-classify yet memorized samples would take more epochs to converge to a final, and equally small, loss value. A similar logic applies to samples with moderately high losses at the end of training. If they are very hard-to-classify samples exhibiting even higher losses on non-member shadow models, they could still be correctly identified by an MIA. Here, a steep drop in loss during training (even if not reaching zero) would signal that a sample is partially memorized.
@@
Fig. 2(a) shows loss traces on three examples from the CIFAR-10 “frog” class: an average example, an easy-to-fit outlier and a hard-to-fit outlier. Loss traces were collected during training, and all samples were part of the training dataset. These traces represent distinct types of loss behavior. The purple sample (average) is a typical sample for the frog class: it is easily learned by the model and the train loss quickly converges to small loss values. The olive sample (easy-to-fit outlier), on the other hand, features unique characteristics compared to other samples from the same class, notably the color. Here we can see that the loss is quite noisy and remains fairly high until approximately epoch 70 when it gets memorized. Finally, the cyan sample (hard-to-fit outlier) is an even harder-to-predict sample from the frog class. The early loss (around epoch 11) for this sample is significantly higher than for the purple example. The loss then decreases, but does not reach the low values exhibited by the purple and olive samples by the end of training. Here, the purple example would be the least vulnerable, and the olive example would be the most vulnerable.
@@
Following this intuition, we want to further verify our hypothesis that easy-to-fit outliers would be the most vulnerable, followed by hard-to-fit outliers, and the average samples would be the least vulnerable. To confirm this, we collect 100 samples from each group and measure the proportion of samples considered vulnerable by LiRA [5] at $\alpha=10^{-3}$ (see Def. 3.1) in these groups. In particular, on Figure 2(b) we consider average loss traces for two groups with 100 samples each, alongside the average train loss. Both groups are selected to have relatively high loss at an early sample in the training (epoch 11, vertical dashed line on the graph) making them more likely to be outliers and at-risk samples. One group then converges to near-zero training loss, representing easy-to-fit outliers. The final loss for the final group remains relatively high, thus representing hard-to-fit outliers. Figure 2(b) indeed confirms that vulnerable samples make up 29% of the easy-to-fit outlier group compared to 12% for hard-to-fit outliers. When comparing to the average baseline of 5%, we confirm that samples with high loss at the early stage of training, outliers, are more vulnerable to MIAs and that, among outliers, easy-to-fit samples (defined by the low final loss) are more vulnerable than hard-to-fit ones. Our method aims to utilize this information to identify at-risk samples.
@@ 5.2 Loss trace aggregation
Note that, where relevant, we use early epoch $s^{*}$ to refer to an epoch at which the average test loss $\mathcal{L}_{\text{test}}^{s^{*}}$ has decreased significantly compared to the initial $\mathcal{L}_{\text{test}}^{0}$, but the train/test gap ($\mathcal{L}_{\text{test}}^{s^{*}}-\mathcal{L}_{\text{train}}^{s^{*}}$) remains low. Intuitively, we consider epoch $s^{*}$ as a point where the model has gained some generalization ability, but has not significantly memorized any samples.
@@
Mean loss (LT-Mean)
@@
Loss Trace $L_{p}$ norms (LT-LP-Norm)
$\phi(x,y)=||\text{trace}(x,y)||_{p}$ (10)
Loss Trace Slope (LT-Slope)
Via least squares, we compute a linear regression on the time steps against the trace, taking the gradient. This describes the rate at which the sample is learned.
$\phi(x,y)=\frac{\Sigma_{s=1}^{S}(\ell_{s}(x,y)-\overline{\ell(x,y)})(s-\overline{s})}{\Sigma_{s=1}^{S}(s-\overline{s})^{2}}$ (11)
Loss Trace Delta (LT-Delta)
@@ 6.1.1 Target model
Adopting the evaluation setup from the literature [5], we train WideResNet28-2 [49] models on the CIFAR-10 dataset [19] for 100 epochs, reaching a test accuracy of $\sim$91%. We train using a stochastic gradient descent optimizer, with a momentum of 0.9 and weight decay set to 0.0005. We use a cosine annealing learning rate scheduler with initial value of 0.1 and a batch size of 256. During training, we augment the images via random horizontal flips, and random crops with four pixels of padding, following the augmentation setup used for training by Carlini et al. [5].
@@
For additional experiments, we vary the model or the dataset used. This involves training WideResNet28-2 models on CIFAR-100 [19] & CINIC10 [10], as well as training ResNet-20 [15] & WideResNet40-4 [49] models on CIFAR-10. In these cases, the training setup is otherwise as previously detailed. In all setups, the train accuracy exceeds 99%.
@@ 6.1.2 Membership Inference Attacks
We evaluate four distinct membership inference attacks: LiRA, Attack R, RMIA and LOSS. We follow the setup proposed in Carlini et al. [5], instantiating online LiRA with 256 shadow models (128 member and 128 non-member models respectively). We instantiate online RMIA with 256 shadow models, and instantiate Attack R with 128 non-member shadow models. All shadow models are trained using the same training procedure described for the target model in Sec. 6.1.1, with each shadow model trained on a different member/non-member split. For LiRA, we use 2 queries per sample through horizontal flipping, matching their “mirror augmentation” setup.
Figure 3: ROC curves for LiRA, Attack R, RMIA and LOSS membership inference attacks on CIFAR-10 against WideResNet28-2 (log scale).
Table 6 shows the performance of MIAs we consider in our training setup. We evaluate them on a random balanced split of members and non-members from the CIFAR-10 train split. As expected, SOTA shadow model-based attacks (LiRA, Attack R, RMIA) significantly outperform LOSS, with LiRA showing the best performance. Note however that though the LOSS attack shows non-trivial AUC=$0.61$, it fails to confidently identify members - as shown by the near-random chance TPR at low FPR values (see Fig. 3).
6.2 Vulnerability to LiRA
@@
Loss 0.21 0.021 0.20 0.060 0.20 0.102
Confidence 0.20 0.020 0.18 0.056 0.20 0.100
Parameter Gradient Norm 0.20 0.020 0.19 0.059 0.21 0.108
Input Gradient Norm 0.20 0.020 0.20 0.060 0.21 0.106
SHAP Values 0.13 0.013 0.14 0.043 0.14 0.071
LT-IQR 0.92 0.094 0.83 0.26 0.76 0.39
Table 1: Precision@k and Recall@k on identifying points vulnerable to LiRA attack. Reported at fixed FPR=$10^{-3}$ and varying values of $k$ for WideResNet28-2 on CIFAR-10
We first compare our method to attack risk predictors from the literature (See Sec. 4.3). We show our method to successfully identify the points most vulnerable to the LiRA MIA and strongly outperform prior methods. Table 1 shows that at a fixed false positive rate of $10^{-3}$, LT-IQR achieves remarkable precision of 92% at $k=1$% (250 samples), substantially outperforming conventional approaches such as loss or gradient norm. This performance advantage is maintained even as we increase $k$ to include a larger portion of training samples, with LT-IQR maintaining 76% precision at k=5% while other methods remain around 20%.
From a recall perspective, LT-IQR demonstrates consistent improvement as k increases, rising from $9\%$ at $k=1\%$ to $39\%$ at $k=5\%$ while maintaining a relatively high precision of $76\%$. This indicates that our method successfully identifies a significant portion of truly vulnerable samples when allowed to make more predictions. In contrast, traditional metrics show minimal recall improvement with increasing $k$, never exceeding $11\%$ even at $k=5\%$.
The stark performance gap between LT-IQR and traditional metrics highlights the importance of considering training dynamics rather than just final model states when assessing sample vulnerability. Single-point metrics like final loss or gradient norms, despite their widespread use in privacy analysis, appear to capture limited information about a sample’s true vulnerability to membership inference attacks. SHAP values, which attempt to quantify feature importance, perform particularly poorly with precision below $0.15$, suggesting that feature attribution alone is insufficient for identifying vulnerable samples.
RN-20 WRN28-2 WRN40-4
Loss 0.21 0.039 0.21 0.021 0.11 0.012
Confidence 0.19 0.035 0.20 0.020 0.10 0.011
Parameter Gradient Norm 0.17 0.031 0.20 0.020 0.12 0.013
Input Gradient Norm 0.19 0.035 0.20 0.020 0.12 0.014
SHAP Values 0.06 0.011 0.13 0.014 0.15 0.017
LT-IQR 0.79 0.145 0.92 0.094 0.91 0.105
Table 2: Precision@k and Recall@k on identifying points vulnerable to LiRA attack. Reported at fixed FPR=$10^{-3}$ and $k=1\%$ for varying model architectures on CIFAR-10
CIFAR-10 CIFAR-100 CINIC-10
Loss 0.21 0.021 0.27 0.011 0.30 0.023
Confidence 0.20 0.021 0.24 0.009 0.28 0.021
Parameter Gradient Norm 0.20 0.020 0.22 0.009 0.29 0.022
Input Gradient Norm 0.20 0.020 0.23 0.009 0.30 0.023
SHAP Values 0.13 0.014 0.18 0.007 0.23 0.018
LT-IQR 0.92 0.094 0.97 0.040 0.94 0.072
Table 3: Precision@k and Recall@k on identifying points vulnerable to LiRA attack. Reported at fixed FPR=$10^{-3}$ and $k=1\%$ for varying datasets using WideResNet28-2
Dataset Model AUC TPR@FPR=$10^{-3}$
CIFAR-10 RN-20 0.70 0.054
WRN40-4 0.73 0.086
WRN28-2 0.74 0.097
CIFAR-100 0.92 0.240
CINIC-10 0.80 0.129
Table 4: LiRA performance on various models and datasets
To validate the robustness of our method, we evaluate LT-IQR across different model architectures and datasets. Note that the underlying LiRA attack shows varying base effectiveness across these configurations (shown in Table 4), resulting in a different number of vulnerable points for each combination of the model and the dataset.
First, we run the evaluation on CIFAR-10 dataset with three model architectures: ResNet-20 (RN-20), WideResNet28-2 (WRN28-2), and WideResNet40-4 (WRN40-4), with results presented in Table 2. Across architectural variants LT-IQR maintains strong performance, achieving $79\%$, $92\%$, and $91\%$ precision respectively at $k=1\%$, while traditional metrics remain around $20\%$.
Second, we run the evaluation with WideResNet28-2 on three datasets: CIFAR-10, CIFAR-100, and CINIC-10 (Table 3). LT-IQR maintains its discriminative power - achieving even higher precision of $97\%$ on CIFAR-100 and $94\%$ on CINIC-10. While recall appears lower on CIFAR-100 ($0.040$ vs $0.094$ on CIFAR-10), this is expected as the stronger attack identifies a larger set of vulnerable points, making it harder to capture a fixed fraction with the same $k$.
We present further results for varying values of $k$ in Appendix A.
Figure 4: Venn diagram comparing the overlap between the set of points predicted to be vulnerable by LiRA, Attack R and RMIA at FPR=$10^{-3}$.
6.3 Cross-attack vulnerability
So far we have only considered vulnerability to one attack, LiRA. It’s worth noting, however, that when evaluating privacy risks, model developers cannot predict which specific membership inference attack an adversary might use. While LiRA generally shows the strongest performance and would be a natural choice for attackers, points that are confidently identified as members by other attacks, such as Attack R or RMIA, also represent significant privacy risks.
To understand the relationships between different attacks’ vulnerabilities, we first analyze how much agreement exists between state-of-the-art attacks in identifying vulnerable points. We consider three widely adopted shadow model-based MIAs: LiRA, Attack R, and RMIA, examining the overlap in points they identify as vulnerable at FPR=$10^{-3}$.
Our analysis reveals that LiRA’s coverage is quite comprehensive: approximately 93% of all points identified as vulnerable by any attack are also identified by LiRA (see Fig. 4). However, there remains a notable 7% of vulnerable points that are only identified by Attack R and/or RMIA. This indicates that while LiRA is highly effective, focusing solely on LiRA vulnerability could miss some genuinely at-risk samples.
Based on this observation, we expand our definition of vulnerable points to include any training sample identified as vulnerable by any of these three attacks. We define this set as the union of vulnerable points across all attacks. As shown in Table 5, our LT-IQR method maintains its strong performance when evaluated against this broader vulnerability definition, achieving Precision@k of $0.93$ for identifying points vulnerable to any attack.
Remarkably, Table 6 shows that when evaluated with this expanded definition of vulnerability, our method’s performance is competitive with shadow model-based approaches themselves. At $k=1\%$, LT-IQR’s $0.93$ precision is comparable to RMIA with 2 shadow models ($0.96$) and outperforms Attack R with 8 models ($0.74$), while requiring only a fraction of their computational cost.
Union
Method Precision@k Recall@k
Loss (desc) 0.21 0.020
Confidence 0.20 0.019
Parameter Gradient Norm 0.20 0.020
Input Gradient Norm 0.20 0.020
SHAP Values 0.17 0.016
LT-IQR 0.93 0.089
Table 5: Precision@k and Recall@k at predicting the vulnerable points defined by the union of three attacks: LiRA, Attack R and RMIA. Reported at fixed FPR=$10^{-3}$ and varying values of $k$ for WideResNet28-2 on CIFAR-10
MIA performance Precision@k
Method # Shadow models AUC TPR@FPR=$10^{-3}$ k=1% k=3% k=5% k=10%
LiRA 256 0.74 0.097 1.00 1.00 1.00 0.98
64 0.73 0.066 1.00 1.00 0.99 0.89
16 0.69 0.046 0.99 0.97 0.92 0.76
Attack R 128 0.66 0.042 1.00 1.00 0.97 0.78
32 0.66 0.000 0.91 0.97 0.94 0.76
8 0.65 0.000 0.74 0.90 0.90 0.75
RMIA 256 0.69 0.050 1.00 1.00 1.00 0.77
64 0.69 0.052 1.00 1.00 0.97 0.75
16 0.67 0.045 1.00 1.00 0.95 0.74
2 0.65 0.026 0.96 0.87 0.80 0.62
LOSS 0.61 0.000 0.11 0.07 0.05 0.05
LT-IQR 0.93 0.83 0.77 0.62
Table 6: Precision@k for traditional MIAs identifying the union of points vulnerable to at least one of the shadow model-based MIAs (LiRA, Attack R, RMIA) with varying number of shadow models. Reported at fixed FPR=$10^{-3}$ on CIFAR-10 with WideResNet28-2)